TETSXPREME:

Mercury is the closest planet to the Sun, with an average distance of 35,000,000 miles. It takes light about 3 minutes to travel from the Sun to Mercury. Neptune is the farthest, with an average distance of 2.8 × 109 miles from the Sun.

TETSXPREME:

Write and simplify an expression showing how many times farther Neptune is from the Sun than Mercury. Write the result in scientific notation.

GalaxyzStarz:

I think it's \[3.5\times 10^7\] For Mercury to Sun And, \[2.8\times10^9\] For Neptune, so I guess you subtract those and you get \[3.5-2.8=0.7\times10^2\] Maybe??

forgetmylife:

@galaxyzstarz wrote:
I think it's \[3.5\times 10^7\] For Mercury to Sun And, \[2.8\times10^9\] For Neptune, so I guess you subtract those and you get \[3.5-2.8=0.7\times10^2\] Maybe??
I thought that as well but I'm still unsure somewhat

TETSXPREME:

Thanks simba i found it

GalaxyzStarz:

;) ofc uncle! :)

TETSXPREME:

8x10